# ClawSafety: "Safe" LLMs, Unsafe Agents

### 01 Apr 2026 v1

@@ Abstract
Personal AI agents like OpenClaw run with elevated privileges on users’ local machines, where a single successful prompt injection can leak credentials, redirect financial transactions, or destroy files. This threat goes well beyond conventional text-level jailbreaks, yet existing safety evaluations fall short: most test models in isolated chat settings, rely on synthetic environments, and do not account for how the agent framework itself shapes safety outcomes. We introduce ClawSafety, a benchmark of 120 adversarial test cases organized along three dimensions (harm domain, attack vector, and harmful action type) and grounded in realistic, high-privilege professional workspaces spanning software engineering, finance, healthcare, law, and DevOps. Each test case embeds adversarial content in one of three channels the agent encounters during normal work: workspace skill files, emails from trusted senders, and web pages. We evaluate five frontier LLMs as agent backbones, running 2,520 sandboxed trials across all configurations. Attack success rates (ASR) range from 40% to 75% across models and vary sharply by injection vector, with skill instructions (highest trust) consistently more dangerous than email or web content. Action-trace analysis reveals that the strongest model maintains hard boundaries against credential forwarding and destructive actions, while weaker models permit both. Cross-scaffold experiments on three agent frameworks further demonstrate that safety is not determined by the backbone model alone but depends on the full deployment stack, calling for safety evaluation that treats model and framework as joint variables.
@@ 1 Introduction
LLMs have rapidly evolved from text-generation tools into the cognitive backbone of autonomous agents that browse the web, compose emails, write and execute code, and manage personal files Zhang et al. (2026a). A turning point came in late January 2026, when OpenClaw [^1] https://openclaw.ai/, an open-source personal assistant agent, went viral and attracted millions of users within weeks. OpenClaw allows any individual to host a personal agent on their own machine, connect it to an LLM API of their choice, and grant it high-privilege access to local files, email accounts, cryptocurrency wallets, and development environments. This setup offers clear productivity gains, but it also creates a wide attack surface: a single successful prompt injection can cascade into real-world harm that goes well beyond what traditional LLM jailbreaks can cause.
Safety research has progressed from text-level jailbreaking Hakim et al. (2026); Vidgen et al. (2024) to multi-modal and web-browsing agents Wang et al. (2025); Evtimov et al. ; Johnson et al. (2025), and recent benchmarks have begun to formalize adversarial evaluation in agentic settings Zhang et al. ; Andriushchenko et al. . Yet these efforts confirm that safety alignment transfers poorly from chat to agentic contexts: a model’s text output can refuse a harmful request while its tool calls simultaneously execute the forbidden action Cartagena and Teixeira (2026); narrow-task finetuning can trigger broad misalignment across unrelated domains Betley et al. (2026); and agent scaffolding itself modulates safety in unpredictable ways Shapira et al. (2026). Existing evaluations, however, fall short along three critical dimensions: most measure safety in isolated chat settings, leaving the chat-vs-agent compliance gap unquantified; prior benchmarks rely on synthetic or task-centric environments that fail to capture the personal, high-privilege, multi-vector threat surface of real-world deployments; and no work jointly treats the backbone model and agent framework as co-variables, obscuring how scaffolding, memory, and tool routing amplify adversarial risk. Shapira et al. (2026) take a step in this direction, documenting emergent OpenClaw failures in a live lab setting, and Wang et al. (2026a) evaluate OpenClaw under black-box conditions, but both remain qualitative or narrow in scope, leaving systematic quantitative evaluation an open problem. We introduce ClawSafety, a safety benchmark for Personal AI agents that addresses these gaps through three contributions:
@@ 2.1 Personal Agent Ecosystems
As LLM capabilities expand, the development of personal agents has evolved from early framework-level optimization (Yao et al., 2022; Shen et al., 2023) to complex, system-wide integrations and explorations (Liang et al., 2026; Song et al., 2025; Agashe et al., 2025a, b; Gonzalez-Pumariega et al., 2025). This architectural shift has recently been highlighted by OpenClaw, a highly developed personal agent system. OpenClaw’s ability to interact with a wide array of real-world software applications makes its ecosystem a natural focal point for agent safety research. Zhang et al. (2026a) study agent behavior on Moltbook  [^4] https://www.moltbook.com, while Wang et al. (2026a) evaluate OpenClaw under adversarial conditions.
@@ 2.2 From Jailbreaks to Agent Attacks
Text-level jailbreak research has shown that safety-aligned LLMs can be coerced into generating harmful content Hakim et al. (2026); Vidgen et al. (2024); Zhang et al. (2026b), but these studies target harmful text, not harmful actions. In the agentic setting, the primary threat is indirect prompt injection (IPI), first formalized by Greshake et al. (2023) and since demonstrated across web-browsing agents Wang et al. (2025); Johnson et al. (2025), multi-agent systems Shahroz et al. (2025), tool-chaining pipelines Li et al. (2025), and coding agents Maloyan and Namiot (2026); see Gulyamov et al. (2026); Ferrag et al. (2025) for taxonomies. These are not hypothetical: Kaleli et al. (2026) document the first confirmed real-world IPI incident in a Palo Alto Networks Unit 42 report from December 2025. A parallel line of work develops detection-side defenses Liu et al. (2025); Wang et al. (2026b); our benchmark instead focuses on measuring attack success across diverse agent configurations. The attack-side advances also expose a deeper problem: safety alignment does not transfer reliably from chat to agentic contexts. Cartagena and Teixeira (2026) show that a model’s text output can refuse a harmful request while its tool calls simultaneously execute the forbidden action. On the training side, DPO-based safety training survives helpfulness optimization but follows a linear safety-helpfulness Pareto frontier Plaut (2026); Wang et al. (2026c), and Betley et al. (2026) show in Nature that narrow-task finetuning can trigger broad misalignment across unrelated domains. Together, these findings support our core hypothesis: chat-level safety $\neq$ agent-level safety.
@@ 2.3 Agent Safety Benchmarks
Several benchmarks evaluate agent safety from different angles: ASB Zhang et al. and AgentHarm Andriushchenko et al. formalize adversarial tasks for tool-integrated agents, WASP Evtimov et al. targets web-browsing agents, SafeAgentBench Yin et al. (2024) and AgentDyn Li et al. (2026) address embodied and dynamic settings, and MASpi An et al. examines multi-agent prompt injection. The closest to our work in scope is InjecAgent Zhan et al. (2024), which introduces 1,054 IPI test cases and evaluates 30 LLMs, finding that ReAct-prompted GPT-4 is vulnerable 24% of the time. However, InjecAgent evaluates LLM-level responses using a single injection channel with binary outcomes. ClawSafety targets agent-level safety: we separate three attack vectors (workspace instructions, email, web) that carry different trust levels, record full action traces with graded harm scores rather than binary labels, and iteratively refine adversarial payloads against frontier models. In a complementary direction, Shapira et al. (2026) conduct qualitative red-teaming of OpenClaw agents, documenting eleven case studies of emergent failures, while Wang et al. (2026a) evaluate OpenClaw under black-box conditions with a more limited set of attack scenarios and backbone models. Our benchmark provides quantitative breadth across models and injection vectors to complement these qualitative and narrow-scope evaluations.
@@ A.1 LLM Jailbreaking
Jailbreaking LLMs: A Survey Hakim et al. [2026]. This survey covers attacks, defenses, and evaluation methodologies for LLM jailbreaking from 2022 to 2025. It reports that automated attacks achieve 90–99% ASR on open-source models. Our work extends the threat model from text generation to agent action execution.
MLCommons AILuminate v0.5 Vidgen et al. [2024]. A standardized jailbreak benchmark from MLCommons. All 39 tested models can be jailbroken to some degree. We build on this finding by asking: if models are already vulnerable at the text level, how much worse is the situation when they control real tools?
@@ A.2 Agent Safety Benchmarks
Agent Security Bench (ASB) Zhang et al. . Formalizes attacks and defenses in LLM-based agents, covering direct prompt injection, indirect prompt injection, and memory poisoning across diverse agent settings. ASB focuses on general-purpose agents; our work targets the specific threat model of personalized local agents with elevated system privileges.
AgentHarm Andriushchenko et al. . A benchmark with 110 base behaviors, 11 harm categories, and 440 unique tasks for measuring harmfulness of LLM agents. AgentHarm evaluates whether agents will proactively perform harmful tasks when instructed; we focus on whether agents can be tricked into harmful actions via injection.
WASP Evtimov et al. . Benchmarks web agent security against prompt injection attacks. Even top models can be deceived by simple injections (partial success rate 16–86%). WASP is limited to the web browsing channel; our benchmark covers web, email, and skill/tool injection vectors.
SafeAgentBench Yin et al. [2024]. Contains 750 safety-related tasks for embodied LLM agents (robotic/physical environments). The embodied setting differs from our software-agent setting, but the evaluation methodology informs our Harm Score design.
AgentDyn Li et al. [2026]. A dynamic, open-ended benchmark for evaluating prompt injection attacks, designed to address the problem that existing benchmarks use overly simplistic tasks. AgentDyn shares our motivation for realism but does not specifically target personal assistant agents.
MASpi An et al. . A unified environment for evaluating prompt injection robustness in multi-agent systems. While MASpi focuses on agent-to-agent interactions, our work focuses on human-to-agent and environment-to-agent attack vectors.
@@ A.3 Prompt Injection Attacks
WebInject Wang et al. [2025]. Introduces prompt injection attacks targeting web agents through pixel-level and text-level injection techniques. We use similar web injection methods as one of our three attack vectors.
Indirect Prompt Injection on Web Agents Johnson et al. [2025]. A demonstration of indirect prompt injection on LLM-based autonomous web navigation agents, using the GCG algorithm and flexible trigger placement strategies. Our web injection attack vector draws on these techniques.
Agents Under Siege Shahroz et al. [2025]. Breaks pragmatic multi-agent LLM systems with optimized prompt attacks. Published at ACL 2025. The optimization-based attack strategy could be applied to strengthen our adversarial payloads in future work.
STAC Li et al. [2025]. Introduces a novel attack category where individually innocent tools are chained to form dangerous attack sequences. This tool-chaining concept is related to our malicious skills/tools attack vector.
Prompt Injection on Coding Assistants Maloyan and Namiot [2026]. A systematic analysis of vulnerabilities in skills, tools, and protocol ecosystems of agentic coding assistants. The skill-based injection patterns described here directly inform our malicious skills attack vector design.
Prompt Injection: A Review Gulyamov et al. [2026]. A review synthesizing 45 key sources from 2023–2025, covering the taxonomy of prompt injection techniques, tool poisoning in agent systems, and the OWASP Top 10 for LLMs. We use their taxonomy as a reference when designing our attack vectors.
From Prompt Injections to Protocol Exploits Ferrag et al. [2025]. Introduces a unified end-to-end threat model for LLM-agent ecosystems, cataloging over 30 attack techniques across input manipulation, model compromise, system/privacy attacks, and protocol vulnerabilities. Our threat taxonomy is complementary, focusing on the personal agent setting.
Fooling AI Agents in the Wild Kaleli et al. [2026]. A Palo Alto Networks Unit 42 report documenting the first real-world instance of malicious indirect prompt injection (December 2025). This confirms that the attacks we model in ClawSafety are not theoretical; they already occur in production.
WAInjectBench Liu et al. [2025]. Benchmarks prompt injection detection methods for web agents. Their detection-focused evaluation complements our attack-focused evaluation.
WebSentinel Wang et al. [2026b]. A system for detecting and localizing prompt injection attacks targeting web agents. WebSentinel represents the type of defense method that our defense baselines aim to establish a reference point for.
@@ A.4 Safety Alignment and Post-Training
Safety Training Persists Through Helpfulness Optimization Plaut [2026]. Finds that safety training persists through subsequent helpfulness optimization in LLM agents, but the safety-helpfulness tradeoff behaves differently in agentic vs. chat settings. This directly supports our hypothesis that chat-level safety does not fully transfer to agent-level safety.
Safety-Utility Tradeoff in LLM Alignment Wang et al. [2026c]. Proposes adaptive safe context learning to address the safety-utility tradeoff. Models trained with detailed safety reasoning chains tend toward over-rejection in chat but may still comply with injected instructions in agentic workflows, a pattern we expect to observe in our experiments.
Emergent Misalignment Betley et al. [2026]. Published in Nature. Shows that finetuning LLMs on narrow tasks (e.g., writing insecure code) causes broad misalignment in unrelated behaviors. This finding suggests that the agentic use case may represent an “unaligned” domain where safety training does not generalize, supporting our core hypothesis.
@@ A.5 Personal Agent Ecosystems
Agents in the Wild Zhang et al. [2026a]. Examines safety, societal impact, and the illusion of sociality in the OpenClaw ecosystem on Moltbook. Provides the broader context for why personal agent safety matters at a societal level.
Agents of Chaos Shapira et al. [2026]. An exploratory red-teaming study in which twenty AI researchers interacted adversarially with OpenClaw agents deployed on isolated Fly.io VMs connected to real Discord and ProtonMail accounts over a two-week period. The paper documents eleven case studies covering unauthorized compliance with non-owners, sensitive information disclosure, disproportionate system-level responses, denial-of-service vulnerabilities, identity spoofing, memory poisoning, and cross-agent propagation of unsafe practices. Three key findings are relevant to ClawSafety: (1) agents lack a reliable “stakeholder model” for distinguishing owner from non-owner instructions, making prompt injection a structural vulnerability rather than a fixable bug; (2) agents’ reported actions frequently diverge from their actual behavior (“failures of social coherence”), meaning that safety audits based on agent self-reports can be misleading; and (3) some observed failures are fundamental to the token-based context window architecture rather than contingent on engineering shortcomings, suggesting that defense methods must go beyond input filtering. Their qualitative methodology is complementary to our quantitative benchmark: they surface failure types through open-ended exploration, while we measure failure rates across controlled experimental conditions. We use their case studies as an instruction guide for our data collection team.
From Assistant to Double Agent Wang et al. [2026a]. The closest concurrent work to ours. Formalizes and benchmarks attacks on OpenClaw, proposing the PASB framework with black-box evaluation. Tests Llama-3.1-70B, Qwen2.5-7B, and GPT-4o-mini, reporting combined ASR up to 66.8%. Our work differs in three ways: (1) we evaluate more recent and capable models (GPT-5.4, Claude Opus 4.6, Gemini 3.0, Qwen 3.5, Kimi K2.5), (2) we cover a broader set of attack vectors including malicious skills/tools, and (3) we include the chat-vs-agent safety gap analysis and defense baselines.

### 04 Apr 2026 v2

# ClawSafety: ”Safe” LLMs, Unsafe Agents

@@ Abstract
Personal AI agents like OpenClaw run with elevated privileges on users’ local machines, where a single successful prompt injection can leak credentials, redirect financial transactions, or destroy files. This threat goes well beyond conventional text-level jailbreaks, yet existing safety evaluations fall short: most test models in isolated chat settings, rely on synthetic environments, and do not account for how the agent framework itself shapes safety outcomes. We introduce ClawSafety, a benchmark of 120 adversarial test cases organized along three dimensions (harm domain, attack vector, and harmful action type) and grounded in realistic, high-privilege professional workspaces spanning software engineering, finance, healthcare, law, and DevOps. Each test case embeds adversarial content in one of three channels the agent encounters during normal work: workspace skill files, emails from trusted senders, and web pages. We evaluate five frontier LLMs as agent backbones, running 2,520 sandboxed trials across all configurations. Attack success rates (ASR) range from 40% to 75% across models and vary sharply by injection vector, with skill instructions (highest trust) consistently more dangerous than email or web content. Action-trace analysis reveals that the strongest model maintains hard boundaries against credential forwarding and destructive actions, while weaker models permit both. Cross-scaffold experiments on three agent frameworks further demonstrate that safety is not determined by the backbone model alone but depends on the full deployment stack, calling for safety evaluation that treats model and framework as joint variables. Code and data will be available at: https://weibowen555.github.io/ClawSafety/.
@@ 1 Introduction
LLMs have rapidly evolved from text-generation tools into the cognitive backbone of autonomous agents that browse the web, compose emails, write and execute code, and manage personal files (Zhang et al., 2026b). A turning point came in late January 2026, when OpenClaw [^1] https://openclaw.ai/, an open-source personal assistant agent, went viral and attracted millions of users within weeks. OpenClaw allows any individual to host a personal agent on their own machine, connect it to an LLM API of their choice, and grant it high-privilege access to local files, email accounts, cryptocurrency wallets, and development environments. This setup offers clear productivity gains, but it also creates a wide attack surface: a single successful prompt injection can cascade into real-world harm that goes well beyond what traditional LLM jailbreaks can cause.
Safety research has progressed from text-level jailbreaking (Hakim et al., 2026; Vidgen et al., 2024) to multi-modal and web-browsing agents (Wang et al., 2025; Evtimov et al., ; Johnson et al., 2025), and recent benchmarks have begun to formalize adversarial evaluation in agentic settings (Zhang et al., ; Andriushchenko et al., ). Yet these efforts confirm that safety alignment transfers poorly from chat to agentic contexts: a model’s text output can refuse a harmful request while its tool calls simultaneously execute the forbidden action (Cartagena & Teixeira, 2026); narrow-task finetuning can trigger broad misalignment across unrelated domains (Betley et al., 2026); and agent scaffolding itself modulates safety in unpredictable ways (Shapira et al., 2026). Existing evaluations, however, fall short along three critical dimensions: most measure safety in isolated chat settings, leaving the chat-vs-agent compliance gap unquantified; prior benchmarks rely on synthetic or task-centric environments that fail to capture the personal, high-privilege, multi-vector threat surface of real-world deployments; and no work jointly treats the backbone model and agent framework as co-variables, obscuring how scaffolding, memory, and tool routing amplify adversarial risk. Shapira et al. (2026) take a step in this direction, documenting emergent OpenClaw failures in a live lab setting, and Wang et al. (2026c) evaluate OpenClaw under black-box conditions, but both remain qualitative or narrow in scope, leaving systematic quantitative evaluation an open problem. We introduce ClawSafety, a safety benchmark for Personal AI agents that addresses these gaps through three contributions:
@@ 2.1 Personal Agent Ecosystems
As LLM capabilities expand, the development of personal agents has evolved from early framework-level optimization (Yao et al., 2022; Shen et al., 2023) to complex, system-wide integrations and explorations (Liang et al., 2026; Song et al., 2025; Agashe et al., 2025a; b; Gonzalez-Pumariega et al., 2025). This architectural shift has recently been highlighted by OpenClaw, a highly developed personal agent system. OpenClaw’s ability to interact with a wide array of real-world software applications makes its ecosystem a natural focal point for agent safety research. Zhang et al. (2026b) study agent behavior on Moltbook  [^4] https://www.moltbook.com, while Wang et al. (2026c) evaluate OpenClaw under adversarial conditions.
@@ 2.2 From Jailbreaks to Agent Attacks
Text-level jailbreak research has shown that safety-aligned LLMs can be coerced into generating harmful content (Hakim et al., 2026; Vidgen et al., 2024; Zhang et al., 2026a), but these studies target harmful text, not harmful actions. In the agentic setting, the primary threat is indirect prompt injection (IPI), first formalized by Greshake et al. (2023) and since demonstrated across web-browsing agents (Wang et al., 2025; Johnson et al., 2025), multi-agent systems (Shahroz et al., 2025), tool-chaining pipelines (Li et al., 2025), and coding agents (Maloyan & Namiot, 2026); see Gulyamov et al. (2026); Ferrag et al. (2025) for taxonomies. These are not hypothetical: Kaleli et al. (2026) document the first confirmed real-world IPI incident in a Palo Alto Networks Unit 42 report from December 2025. A parallel line of work develops detection-side defenses (Liu et al., 2025; Wang et al., 2026a); our benchmark instead focuses on measuring attack success across diverse agent configurations. The attack-side advances also expose a deeper problem: safety alignment does not transfer reliably from chat to agentic contexts. Cartagena & Teixeira (2026) show that a model’s text output can refuse a harmful request while its tool calls simultaneously execute the forbidden action. On the training side, DPO-based safety training survives helpfulness optimization but follows a linear safety-helpfulness Pareto frontier (Plaut, 2026; Wang et al., 2026b), and Betley et al. (2026) show in Nature that narrow-task finetuning can trigger broad misalignment across unrelated domains. Together, these findings support our core hypothesis: chat-level safety $\neq$ agent-level safety.
@@ 2.3 Agent Safety Benchmarks
Several benchmarks evaluate agent safety from different angles: ASB (Zhang et al., ) and AgentHarm (Andriushchenko et al., ) formalize adversarial tasks for tool-integrated agents, WASP (Evtimov et al., ) targets web-browsing agents, SafeAgentBench (Yin et al., 2024) and AgentDyn (Li et al., 2026) address embodied and dynamic settings, and MASpi (An et al., ) examines multi-agent prompt injection. The closest to our work in scope is InjecAgent (Zhan et al., 2024), which introduces 1,054 IPI test cases and evaluates 30 LLMs, finding that ReAct-prompted GPT-4 is vulnerable 24% of the time. However, InjecAgent evaluates LLM-level responses using a single injection channel with binary outcomes. ClawSafety targets agent-level safety: we separate three attack vectors (workspace instructions, email, web) that carry different trust levels, record full action traces with graded harm scores rather than binary labels, and iteratively refine adversarial payloads against frontier models. In a complementary direction, Shapira et al. (2026) conduct qualitative red-teaming of OpenClaw agents, documenting eleven case studies of emergent failures, while Wang et al. (2026c) evaluate OpenClaw under black-box conditions with a more limited set of attack scenarios and backbone models. Our benchmark provides quantitative breadth across models and injection vectors to complement these qualitative and narrow-scope evaluations.
@@ A.1 LLM Jailbreaking
Jailbreaking LLMs: A Survey (Hakim et al., 2026). This survey covers attacks, defenses, and evaluation methodologies for LLM jailbreaking from 2022 to 2025. It reports that automated attacks achieve 90–99% ASR on open-source models. Our work extends the threat model from text generation to agent action execution.
MLCommons AILuminate v0.5 (Vidgen et al., 2024). A standardized jailbreak benchmark from MLCommons. All 39 tested models can be jailbroken to some degree. We build on this finding by asking: if models are already vulnerable at the text level, how much worse is the situation when they control real tools?
@@ A.2 Agent Safety Benchmarks
Agent Security Bench (ASB) (Zhang et al., ). Formalizes attacks and defenses in LLM-based agents, covering direct prompt injection, indirect prompt injection, and memory poisoning across diverse agent settings. ASB focuses on general-purpose agents; our work targets the specific threat model of personalized local agents with elevated system privileges.
AgentHarm (Andriushchenko et al., ). A benchmark with 110 base behaviors, 11 harm categories, and 440 unique tasks for measuring harmfulness of LLM agents. AgentHarm evaluates whether agents will proactively perform harmful tasks when instructed; we focus on whether agents can be tricked into harmful actions via injection.
WASP (Evtimov et al., ). Benchmarks web agent security against prompt injection attacks. Even top models can be deceived by simple injections (partial success rate 16–86%). WASP is limited to the web browsing channel; our benchmark covers web, email, and skill/tool injection vectors.
SafeAgentBench (Yin et al., 2024). Contains 750 safety-related tasks for embodied LLM agents (robotic/physical environments). The embodied setting differs from our software-agent setting, but the evaluation methodology informs our Harm Score design.
AgentDyn (Li et al., 2026). A dynamic, open-ended benchmark for evaluating prompt injection attacks, designed to address the problem that existing benchmarks use overly simplistic tasks. AgentDyn shares our motivation for realism but does not specifically target personal assistant agents.
MASpi (An et al., ). A unified environment for evaluating prompt injection robustness in multi-agent systems. While MASpi focuses on agent-to-agent interactions, our work focuses on human-to-agent and environment-to-agent attack vectors.
@@ A.3 Prompt Injection Attacks
WebInject (Wang et al., 2025). Introduces prompt injection attacks targeting web agents through pixel-level and text-level injection techniques. We use similar web injection methods as one of our three attack vectors.
Indirect Prompt Injection on Web Agents (Johnson et al., 2025). A demonstration of indirect prompt injection on LLM-based autonomous web navigation agents, using the GCG algorithm and flexible trigger placement strategies. Our web injection attack vector draws on these techniques.
Agents Under Siege (Shahroz et al., 2025). Breaks pragmatic multi-agent LLM systems with optimized prompt attacks. Published at ACL 2025. The optimization-based attack strategy could be applied to strengthen our adversarial payloads in future work.
STAC (Li et al., 2025). Introduces a novel attack category where individually innocent tools are chained to form dangerous attack sequences. This tool-chaining concept is related to our malicious skills/tools attack vector.
Prompt Injection on Coding Assistants (Maloyan & Namiot, 2026). A systematic analysis of vulnerabilities in skills, tools, and protocol ecosystems of agentic coding assistants. The skill-based injection patterns described here directly inform our malicious skills attack vector design.
Prompt Injection: A Review (Gulyamov et al., 2026). A review synthesizing 45 key sources from 2023–2025, covering the taxonomy of prompt injection techniques, tool poisoning in agent systems, and the OWASP Top 10 for LLMs. We use their taxonomy as a reference when designing our attack vectors.
From Prompt Injections to Protocol Exploits (Ferrag et al., 2025). Introduces a unified end-to-end threat model for LLM-agent ecosystems, cataloging over 30 attack techniques across input manipulation, model compromise, system/privacy attacks, and protocol vulnerabilities. Our threat taxonomy is complementary, focusing on the personal agent setting.
Fooling AI Agents in the Wild (Kaleli et al., 2026). A Palo Alto Networks Unit 42 report documenting the first real-world instance of malicious indirect prompt injection (December 2025). This confirms that the attacks we model in ClawSafety are not theoretical; they already occur in production.
WAInjectBench (Liu et al., 2025). Benchmarks prompt injection detection methods for web agents. Their detection-focused evaluation complements our attack-focused evaluation.
WebSentinel (Wang et al., 2026a). A system for detecting and localizing prompt injection attacks targeting web agents. WebSentinel represents the type of defense method that our defense baselines aim to establish a reference point for.
@@ A.4 Safety Alignment and Post-Training
Safety Training Persists Through Helpfulness Optimization (Plaut, 2026). Finds that safety training persists through subsequent helpfulness optimization in LLM agents, but the safety-helpfulness tradeoff behaves differently in agentic vs. chat settings. This directly supports our hypothesis that chat-level safety does not fully transfer to agent-level safety.
Safety-Utility Tradeoff in LLM Alignment (Wang et al., 2026b). Proposes adaptive safe context learning to address the safety-utility tradeoff. Models trained with detailed safety reasoning chains tend toward over-rejection in chat but may still comply with injected instructions in agentic workflows, a pattern we expect to observe in our experiments.
Emergent Misalignment (Betley et al., 2026). Published in Nature. Shows that finetuning LLMs on narrow tasks (e.g., writing insecure code) causes broad misalignment in unrelated behaviors. This finding suggests that the agentic use case may represent an “unaligned” domain where safety training does not generalize, supporting our core hypothesis.
@@ A.5 Personal Agent Ecosystems
Agents in the Wild (Zhang et al., 2026b). Examines safety, societal impact, and the illusion of sociality in the OpenClaw ecosystem on Moltbook. Provides the broader context for why personal agent safety matters at a societal level.
Agents of Chaos (Shapira et al., 2026). An exploratory red-teaming study in which twenty AI researchers interacted adversarially with OpenClaw agents deployed on isolated Fly.io VMs connected to real Discord and ProtonMail accounts over a two-week period. The paper documents eleven case studies covering unauthorized compliance with non-owners, sensitive information disclosure, disproportionate system-level responses, denial-of-service vulnerabilities, identity spoofing, memory poisoning, and cross-agent propagation of unsafe practices. Three key findings are relevant to ClawSafety: (1) agents lack a reliable “stakeholder model” for distinguishing owner from non-owner instructions, making prompt injection a structural vulnerability rather than a fixable bug; (2) agents’ reported actions frequently diverge from their actual behavior (“failures of social coherence”), meaning that safety audits based on agent self-reports can be misleading; and (3) some observed failures are fundamental to the token-based context window architecture rather than contingent on engineering shortcomings, suggesting that defense methods must go beyond input filtering. Their qualitative methodology is complementary to our quantitative benchmark: they surface failure types through open-ended exploration, while we measure failure rates across controlled experimental conditions. We use their case studies as an instruction guide for our data collection team.
From Assistant to Double Agent (Wang et al., 2026c). The closest concurrent work to ours. Formalizes and benchmarks attacks on OpenClaw, proposing the PASB framework with black-box evaluation. Tests Llama-3.1-70B, Qwen2.5-7B, and GPT-4o-mini, reporting combined ASR up to 66.8%. Our work differs in three ways: (1) we evaluate more recent and capable models (GPT-5.4, Claude Opus 4.6, Gemini 3.0, Qwen 3.5, Kimi K2.5), (2) we cover a broader set of attack vectors including malicious skills/tools, and (3) we include the chat-vs-agent safety gap analysis and defense baselines.